The present value of an annuity is R62 543,42. The time under Consideration is 10 years and the applicable interest rate is 16% per year, compounded monthly. The future value of this annuity is

1. R162 612,89.

2. R1 333 517,61.

3. R306 521,61.

4. R581 913,00.


1
Expert's answer
2022-03-31T06:43:32-0400

"F = P * ([1 + I]^{N - 1} )/I\\

F=62546.42(1+0.16)^{12-1}/0.16\\ F=306521.61\\ The \:correct\:answer\:is\:306521.61\\F=62546.42\frac{(1+0.16)^{11}}
 {0.16}\\
F=306521.61"